Physicshelp02:

A shuffleboard disc of mass 0.50 kg accelerates under an applied force of 12.0 N [forward]. if the disk moves from rest for 0.20s, how far does it travel while accelerating? And also what would the final velocity be?

justjm:

First you must find acceleration \(F=ma\) F=12.0 N m=0.5 kg hence, a=12/0.5 = 24.0 m/s^2 Now, I'd like to solve for the second question first, which is finding final velocity. \(v=\int_{ }^{ }a\ dt\) \(v=\int_{ }^{ }24\ dt\ =\ 24t+v_{o}\) Since the disk moves from REST, the initial velocity is 0, and hence the velocity (final velocity) can be represented as \(v=24t\). It moves for 0.20 seconds, so plugging that in, \(0.20\cdot24\) = \(4.8~m/s\) Now back to the first question: how far does it travel while accelerating: \(x=\int_{ }^{ }\left|v\left(t\right)\left|dt=\int_{ }^{ }\right|24t\right|dt=12t^{2}\) t=0.2 hence distance is \(0.2^{2}\cdot12=\) You can calculate that

justjm:

You can also use kinematics equations here for question 2 - \(a\approx\frac{v_{f}-v_{i}}{t}\) For question 1 - \(x=x_{i}+v_{i}t+\frac{1}{2}at^{2}\)